# Stable cohomology of universal character varieties

### v1

@@ 3.2. Smoothness and purity for $d$ coprime to $n$
When $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, the Higgs moduli space $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ is a smooth algebraic variety [Nit91, Proposition 7.4], and the corresponding fine moduli stack is a $\mathbb{C}^{*}$-gerbe over $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$. In this case, $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is also smooth.
@@
The Hitchin morphism $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\to\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$ takes a Higgs bundle $(E,\theta)$ to the coefficients of the characteristic polynomial of $\theta$, viewed as a twisted endomorphism of $E$. The Hitchin morphism is proper [Sim94, Theorem 6.11]. It is also equivariant with respect to the $\mathbb{G}_{m}$-action $t\cdot(E,\phi):=(E,t\phi)$ on the source and $t\cdot x\in H^{0}(C,\omega_{C}^{\otimes i}):=t^{i}x$ on the target. The Hitchin morphism is also invariant under the action of $\mathrm{Jac}(C)[n]$. It therefore factors through a proper morphism $p\colon M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)\to\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$. The $\mathbb{G}_{m}$-action likewise descends to an action on $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$ and, by the properness of $p$, any morphism $\mathbb{G}_{m}\to M_{\mathrm{Dol}}^{\operatorname{PGL}}(C,n,d),t\mapsto t\cdot x$ can be extended to a morphism from $\mathbb{A}^{1}$. The $\mathbb{G}_{m}$-fixed locus $M_{\mathrm{Dol}}^{\operatorname{PGL}}(C,n,d)^{\mathbb{G}_{m}}$ is proper, since it is a closed subspace of the fiber $p^{-1}(0)$.
@@
The proof is essentially identical to that of [HRV15, Corollary 1.3.2]. ∎
@@
The character variety $M_{B}^{\operatorname{PGL}}(C,n,d)$ is homeomorphic to $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$, by the nonabelian Hodge correspondence, but this homeomorphism is not algebraic and the induced isomorphism on cohomology does not preserve weights. Indeed, the weight filtration on $H^{k}(M_{B}^{\operatorname{PGL}}(C,n,d);\mathbb{Q})$ is typically far from pure, and is the subject of the $P=W$ conjecture proposed in [dCHM12]. The $P=W$ conjecture for $\operatorname{PGL}_{n}$ follows from the corresponding conjecture for $\operatorname{GL}_{n}$, which is proved in [MS24, HMMS25].
@@ 3.3. Universal Higgs moduli spaces
$(L,s)\cdot(E,\theta)=(L\otimes E,s+\theta).$
Then (10) also generalizes to the relative setting. Thus, from the universal curve over $\mathcal{M}_{g}$, we obtain a universal $\operatorname{PGL}$-Higgs moduli space $M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}.$ By [Sim94, Theorem 9.11 and Lemma 9.14], there is a homeomorphism from the relative $\operatorname{PGL}$-character variety to the relative $\operatorname{PGL}$-Higgs moduli space for an arbitrary family of curves. Simpson’s homeomorphism preserves the $C^{\infty}$-isomorphism class of the underlying $\operatorname{PGL}$-bundles, and hence maps degree $d$ character varieties to degree $d$ Higgs moduli spaces.
Universal Nonabelian Hodge Correspondence.
@@
Let $\pi_{B}\colon\mathcal{M}^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ be the universal degree $d$ $\operatorname{PGL}$-character variety. Then the pushforward $R\pi_{B*}\mathbb{Q}$ for splits as a sum of its cohomology sheaves if and only if the same is true for $\pi_{\mathrm{Dol}}\colon\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}.$
@@
The universal Higgs moduli space $\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ has a relative Hitchin morphism to the vector bundle whose fiber over $[C]$ is $\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$. Furthermore, there is a natural relative compactification of the universal Higgs moduli space, whose construction is described in [dC21], expanding on [Sim97, §11]. We briefly summarize its essential properties. Define a $\mathbb{G}_{m}$-action on $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\times\mathbb{A}^{1}$ by the standard scaling action on the first factor and $t\cdot x:=tx$ on the second. Let $M^{\circ}\subset M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ be the subset of points such that the $\mathbb{G}_{m}$-action does not have a limit at $\infty$. The quotient $M^{\circ}/\mathbb{G}_{m}]$ is a relative projective compactification of $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ over $\mathcal{M}_{g}$. Moreover, both $M^{\circ}/\mathbb{G}_{m}]$ and the boundary divisor are smooth over $\mathcal{M}_{g}$.
@@
To handle constructible sheaves equipped with Hodge structures, we use Saito’s theory of mixed Hodge modules [Sai90]. The key features are as follows. There is a category of mixed Hodge modules $\operatorname{MHM}(X)$ on each complex variety $X$, with a standard six-functor formalism. There is a functor
@@
taking a mixed Hodge module to its underlying rational structure, which is compatible with the six functors [Sai90, Theorem 0.1]. Moreover, $\operatorname{MHM}(\mathrm{pt})$ is the category of polarizable $\mathbb{Q}$-mixed Hodge structures [Sai90, (4.2.12)], and pushing forward the trivial Hodge module $\mathbb{Q}^{H}_{X}$ to a point we recover Deligne’s mixed Hodge structure on $H^{*}(X;\mathbb{Q})$ [Sai90, p. 328].
@@
Deligne’s theory of mixed Hodge structures extends to Deligne–Mumford stacks and, more generally, simplicial objects in the category of algebraic varieties [Del74]. Saito’s theory likewise extends to Deligne–Mumford stacks, as explained in [Tub25]. We use the basic properties of mixed Hodge structures and mixed Hodge modules on Deligne–Mumford stacks without further comment.
@@
Our goal is to show that the (derived) pushforward $\pi_{*}\mathbb{Q}^{H}_{U}\in D^{b}\operatorname{MHM}(S)$ is pure of weight 0 [Sai90, §4.5], i.e., that $\mathrm{Gr}^{W}_{i}\mathcal{H}^{j}(\pi_{*}\mathbb{Q}^{H}_{U})=0$ for $i\neq j$. This condition guarantees that $\pi_{*}\mathbb{Q}^{H}_{U}\cong\oplus_{i}\mathcal{H}^{i}(\pi_{*}\mathbb{Q}^{H}_{U})[-i]\in D^{b}\operatorname{MHM}(S)$ [Sai90, (4.5.4)], and then applying the rat functor (11), one recovers the same identity on the level of constructible sheaves.
To prove the purity condition, it suffices to check that for every $s\in S$ and $k\in\mathbb{Z}$ we have that $\mathcal{H}^{k}i_{s}^{*}\pi_{*}\mathbb{Q}^{H}_{U}$ is of weight $\leq k$ and $\mathcal{H}^{k}i_{s}^{!}\pi_{*}\mathbb{Q}_{U}^{H}$ is of weight $\geq k$ [Sai90, (4.6.1)]. We start with the identity [Sai90, (4.4.3)]
@@
For the first assertion, for $d:=\dim S$, note that $i_{s}^{*}=i_{s}^{!}(-d)[-2d]$ by smoothness of $S$ and $i_{U_{s}}^{*}=i_{U_{s}}^{!}(-d)[-2d]$ by smoothness of $S$ and $f$ [Sai90, (4.4.1)], and thus
@@
By Corollary 3.10 it suffices to prove the statement on the Dolbeault side, for $\pi\colon\mathcal{M}_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$. Recall that $\pi$ is a smooth morphism [Nit91, Proposition 7.4] whose fibers have pure cohomology, by Proposition 3.8. The theorem then follows from Lemma 3.14.
@@
in $D^{b}_{c}(\mathbb{Q}_{X})$ [Dim04, p. 45], where $j^{-1}\mathbb{Q}_{X}=\mathbb{Q}_{U}$. Since $X$ and $Z$ are smooth, and $Z$ is of complex codimension 1, we have $i^{!}\mathbb{Q}_{X}=\mathbb{Q}_{Z}[-2]$ [Dim04, Proposition 3.2.11]. Pushing forward along the map $\pi\colon X\to\mathcal{M}_{g}$, we obtain an exact triangle
@@
whose long exact sequence of cohomology is the relative Gysin sequence. Restricting to the fiber over any point $s\in\mathcal{M}_{g}$, using proper base change [Dim04, Theorem 2.3.26] to identify $(R\pi_{X*}\mathbb{Q}_{X})_{s}\cong H^{*}(X_{s},\mathbb{Q})$ and similarly for $Z$, and taking the long exact sequence of cohomology, we obtain the Gysin sequence for the smooth divisor $Z_{s}\subset X_{s}$
@@
This is an exact sequence of mixed Hodge structures [Del74, (9.2.1.2)]. By Proposition 3.8, $H^{i}(U_{s})$ is pure of weight $i$. The same is true for the cohomology groups of $X_{s}$ and $Z_{s}$, since both of these spaces are smooth and proper. In particular, this implies that the long exact sequence above breaks into short exact sequences. Now, since this is true on each fiber and vanishing of a map of sheaves can be checked on fibers, we conclude that the relative Gysin sequence splits into short exact sequences
@@
Note that $R^{i}\pi_{*}\mathbb{Q}_{X}$ is a semisimple local system, by [Del68, Proposition 2.1] and [Del71, Cor. 4.2.9(a)]. The same being true for $\pi_{Z}$, we conclude that each $R^{i}\pi_{U*}\mathbb{Q}_{U}$ is also a local system, and each short exact sequence above is split, by semisimplicity of the middle factor.
Theorem 3.1 therefore follows from Lemma 3.15 below, applied to the standard $t$-structure on $D^{b}_{c}(\mathbb{Q}_{\mathcal{M}_{g}})$ and the triangle in equation (12) (cf. [Dim04, Example 5.1.3]). ∎
@@
satisfying the hypotheses of the lemma. To prove this, we apply the 9-lemma [BBD82, Proposition 1.1.11] to the commutative square coming from applying $id\to\tau^{\geq 0}$ to $A\to B$:
@@
First, recall that the splitting in Theorem 3.1 is stated for the universal $\operatorname{PGL}$-character variety over $\mathcal{M}_{g}$. By smooth base change [Dim04, Theorem 3.2.13(ii)], the analogous splitting holds for any family of curves induced by a smooth morphism to $\mathcal{M}_{g}$. In particular, this holds for the universal curve over $\mathcal{M}_{g,1}$.
@@
It remains to prove the claim. By [HP12, Theorem 1.4], we have
@@
All of the functors in this proof are derived; for simplicity we omit this from the notation. We start with the case that $S$ is a manifold. The strategy is to apply Verdier duality to the proper relative Künneth formula [KS90, Exercise II.18(i)]
@@
[Dim04, Theorem 3.2.17(iii) and Proposition 3.3.7]. On the right side we have
@@
where for the first equality we use [KS90, Proposition 2.6.3(ii)]. Identifying $D\mathbb{Q}=\mathbb{Q}[\dim X_{2}]$ and using the assumption that the sheaves $R^{j}\pi_{2*}\mathbb{Q}$ are locally trivial we see from [dCM09, p. 624] that
@@
To show the splitting of $R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}$, we use Simpson’s ($d=0$) nonabelian Hodge theorem $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)\cong M_{\mathrm{Dol}}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)$ [Sim94, Proposition 6.8 and Theorem 7.18] and the fact that, as noted in Remark 3.12, the proof of Theorem 3.1 holds for $M_{\mathrm{Dol}}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)$. Furthermore, $\Gamma$ acts trivially on the cohomology of the fibers. This is because the action on the fibers can be extended to the multiplication action of the connected group $\mathrm{Hom}(\pi_{1}(S_{g,1}),\mathbb{C}^{*})\cong(\mathbb{C}^{*})^{2g}$ on itself. Then, since $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)\to\mathcal{M}_{g,1}$ is topologically locally trivial, we have $(R\pi_{\operatorname{GL}_{1}*}\mathbb{Q})^{\widetilde{\Gamma}}=R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}$. Thus, we conclude that

### v2

@@ 3.2. Smoothness and purity for $d$ coprime to $n$
When $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, the Higgs moduli space $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ is a smooth algebraic variety [41, Proposition 7.4], and the corresponding fine moduli stack is a $\mathbb{C}^{*}$-gerbe over $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$. In this case, $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is also smooth.
@@
The Hitchin morphism $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\to\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$ takes a Higgs bundle $(E,\theta)$ to the coefficients of the characteristic polynomial of $\theta$, viewed as a twisted endomorphism of $E$. The Hitchin morphism is proper [49, Theorem 6.11]. It is also equivariant with respect to the $\mathbb{G}_{m}$-action $t\cdot(E,\phi):=(E,t\phi)$ on the source and $t\cdot x\in H^{0}(C,\omega_{C}^{\otimes i}):=t^{i}x$ on the target. The Hitchin morphism is also invariant under the action of $\mathrm{Jac}(C)[n]$. It therefore factors through a proper morphism $p\colon M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)\to\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$. The $\mathbb{G}_{m}$-action likewise descends to an action on $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$ and, by the properness of $p$, any morphism $\mathbb{G}_{m}\to M_{\mathrm{Dol}}^{\operatorname{PGL}}(C,n,d),t\mapsto t\cdot x$ can be extended to a morphism from $\mathbb{A}^{1}$. The $\mathbb{G}_{m}$-fixed locus $M_{\mathrm{Dol}}^{\operatorname{PGL}}(C,n,d)^{\mathbb{G}_{m}}$ is proper, since it is a closed subspace of the fiber $p^{-1}(0)$.
@@
The proof is essentially identical to that of [18, Corollary 1.3.2]. ∎
@@
The character variety $M_{B}^{\operatorname{PGL}}(C,n,d)$ is homeomorphic to $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$, by the nonabelian Hodge correspondence, but this homeomorphism is not algebraic and the induced isomorphism on cohomology does not preserve weights. Indeed, the weight filtration on $H^{k}(M_{B}^{\operatorname{PGL}}(C,n,d);\mathbb{Q})$ is typically far from pure, and is the subject of the $P=W$ conjecture proposed in [5]. The $P=W$ conjecture for $\operatorname{PGL}_{n}$ follows from the corresponding conjecture for $\operatorname{GL}_{n}$, which is proved in [36, 16].
@@ 3.3. Universal Higgs moduli spaces
$(M,s)\cdot(E,\theta)=(M\otimes E,s+\theta).$
Then (10) also generalizes to the relative setting. Thus, from the universal curve over $\mathcal{M}_{g}$, we obtain a universal $\operatorname{PGL}$-Higgs moduli space $M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}.$ By [49, Theorem 9.11 and Lemma 9.14], there is a homeomorphism from the relative $\operatorname{PGL}$-character variety to the relative $\operatorname{PGL}$-Higgs moduli space for an arbitrary family of curves. Simpson’s homeomorphism preserves the $C^{\infty}$-isomorphism class of the underlying $\operatorname{PGL}$-bundles, and hence maps degree $d$ character varieties to degree $d$ Higgs moduli spaces.
Universal $\operatorname{PGL}$ Nonabelian Hodge Correspondence.
@@
Let $\pi_{B}\colon\mathcal{M}^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ be the universal degree $d$ $\operatorname{PGL}$-character variety. Then the pushforward $R\pi_{B*}\mathbb{Q}$ splits as a sum of its cohomology sheaves if and only if the same is true for $\pi_{\mathrm{Dol}}\colon\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}.$
@@
The universal Higgs moduli space $\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ has a relative Hitchin morphism to the vector bundle whose fiber over $[C]$ is $\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$. Furthermore, there is a natural relative compactification of the universal Higgs moduli space, whose construction is described in [8], expanding on [50, §11]. We briefly summarize its essential properties. Define a $\mathbb{G}_{m}$-action on $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\times\mathbb{A}^{1}$ by the standard scaling action on the first factor and $t\cdot x:=tx$ on the second. Let $M^{\circ}\subset M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ be the subset of points such that the $\mathbb{G}_{m}$-action does not have a limit at $\infty$. The quotient $M^{\circ}/\mathbb{G}_{m}]$ is a relative projective compactification of $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ over $\mathcal{M}_{g}$. Moreover, both $M^{\circ}/\mathbb{G}_{m}]$ and the boundary divisor are smooth over $\mathcal{M}_{g}$.
@@
To handle constructible sheaves equipped with Hodge structures, we use Saito’s theory of mixed Hodge modules [45]. The key features are as follows. There is a category of mixed Hodge modules $\operatorname{MHM}(X)$ on each complex variety $X$, with a standard six-functor formalism. There is a functor
@@
taking a mixed Hodge module to its underlying rational structure, which is compatible with the six functors [45, Theorem 0.1]. Moreover, $\operatorname{MHM}(\mathrm{pt})$ is the category of polarizable $\mathbb{Q}$-mixed Hodge structures [45, (4.2.12)], and pushing forward the trivial Hodge module $\mathbb{Q}^{H}_{X}$ to a point we recover Deligne’s mixed Hodge structure on $H^{*}(X;\mathbb{Q})$ [45, p. 328].
@@
Deligne’s theory of mixed Hodge structures extends to Deligne–Mumford stacks and, more generally, simplicial objects in the category of algebraic varieties [11]. Saito’s theory likewise extends to Deligne–Mumford stacks, as explained in [51]. We use the basic properties of mixed Hodge structures and mixed Hodge modules on Deligne–Mumford stacks without further comment.
@@
Our goal is to show that the (derived) pushforward $\pi_{*}\mathbb{Q}^{H}_{U}\in D^{b}\operatorname{MHM}(S)$ is pure of weight 0 [45, §4.5], i.e., that $\mathrm{Gr}^{W}_{i}\mathcal{H}^{j}(\pi_{*}\mathbb{Q}^{H}_{U})=0$ for $i\neq j$. This condition guarantees that $\pi_{*}\mathbb{Q}^{H}_{U}\cong\oplus_{i}\mathcal{H}^{i}(\pi_{*}\mathbb{Q}^{H}_{U})[-i]\in D^{b}\operatorname{MHM}(S)$ [45, (4.5.4)], and then applying the rat functor (11), one recovers the same identity on the level of constructible sheaves.
To prove the purity condition, it suffices to check that for every $s\in S$ and $k\in\mathbb{Z}$ we have that $\mathcal{H}^{k}i_{s}^{*}\pi_{*}\mathbb{Q}^{H}_{U}$ is of weight $\leq k$ and $\mathcal{H}^{k}i_{s}^{!}\pi_{*}\mathbb{Q}_{U}^{H}$ is of weight $\geq k$ [45, (4.6.1)]. We start with the identity [45, (4.4.3)]
@@
For the first assertion, for $d:=\dim S$, note that $i_{s}^{*}=i_{s}^{!}(-d)[-2d]$ by smoothness of $S$ and $i_{U_{s}}^{*}=i_{U_{s}}^{!}(-d)[-2d]$ by smoothness of $S$ and $f$ [45, (4.4.1)], and thus
@@
By Corollary 3.10 it suffices to prove the statement on the Dolbeault side, for $\pi\colon\mathcal{M}_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$. Recall that $\pi$ is a smooth morphism [41, Proposition 7.4] whose fibers have pure cohomology, by Proposition 3.8. The theorem then follows from Lemma 3.14.
@@
in $D^{b}_{c}(\mathbb{Q}_{X})$ [13, p. 45], where $j^{-1}\mathbb{Q}_{X}=\mathbb{Q}_{U}$. Since $X$ and $Z$ are smooth, and $Z$ is of complex codimension 1, we have $i^{!}\mathbb{Q}_{X}=\mathbb{Q}_{Z}[-2]$ [13, Proposition 3.2.11]. Pushing forward along the map $\pi\colon X\to\mathcal{M}_{g}$, we obtain an exact triangle
@@
whose long exact sequence of cohomology is the relative Gysin sequence. Restricting to the fiber over any point $s\in\mathcal{M}_{g}$, using proper base change [13, Theorem 2.3.26] to identify $(R\pi_{X*}\mathbb{Q}_{X})_{s}\cong H^{*}(X_{s},\mathbb{Q})$ and similarly for $Z$, and taking the long exact sequence of cohomology, we obtain the Gysin sequence for the smooth divisor $Z_{s}\subset X_{s}$
@@
This is an exact sequence of mixed Hodge structures [11, (9.2.1.2)]. By Proposition 3.8, $H^{i}(U_{s})$ is pure of weight $i$. The same is true for the cohomology groups of $X_{s}$ and $Z_{s}$, since both of these spaces are smooth and proper. In particular, this implies that the long exact sequence above breaks into short exact sequences. Now, since this is true on each fiber and vanishing of a map of sheaves can be checked on fibers, we conclude that the relative Gysin sequence splits into short exact sequences
@@
Note that $R^{i}\pi_{*}\mathbb{Q}_{X}$ is a semisimple local system, by [9, Proposition 2.1] and [10, Cor. 4.2.9(a)]. The same being true for $\pi_{Z}$, we conclude that each $R^{i}\pi_{U*}\mathbb{Q}_{U}$ is also a local system, and each short exact sequence above is split, by semisimplicity of the middle factor.
Theorem 3.1 therefore follows from Lemma 3.15 below, applied to the standard $t$-structure on $D^{b}_{c}(\mathbb{Q}_{\mathcal{M}_{g}})$ and the triangle in equation (12) (cf. [13, Example 5.1.3]). ∎
@@
satisfying the hypotheses of the lemma. To prove this, we apply the 9-lemma [2, Proposition 1.1.11] to the commutative square coming from applying $id\to\tau^{\geq 0}$ to $A\to B$:
@@
First, recall that the splitting in Theorem 3.1 is stated for the universal $\operatorname{PGL}$-character variety over $\mathcal{M}_{g}$. By smooth base change [13, Theorem 3.2.13(ii)], the analogous splitting holds for any family of curves induced by a smooth morphism to $\mathcal{M}_{g}$. In particular, this holds for the universal curve over $\mathcal{M}_{g,1}$.
@@
It remains to prove the claim. By [17, Theorem 1.4], we have
@@
All of the functors in this proof are derived; for simplicity we omit this from the notation. Also, in this proof (only) we use $\dim$ to denote $\dim_{\mathbb{R}}$. We start with the case that $S$ is a manifold. The strategy is to apply Verdier duality to the proper relative Künneth formula [25, Exercise II.18(i)]
@@
[13, Theorem 3.2.17(iii) and Proposition 3.3.7]. On the right side we have
@@
where for the first equality we use [25, Proposition 2.6.3(ii)]. Identifying $D\mathbb{Q}=\mathbb{Q}[\dim X_{2}]$ and using the assumption that the sheaves $R^{j}\pi_{2*}\mathbb{Q}$ are locally trivial we see from [7, p. 624] that
@@
To show the splitting of $R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}$, we use Simpson’s ($d=0$) nonabelian Hodge theorem $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)\cong M_{\mathrm{Dol}}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)$ [49, Proposition 6.8 and Theorem 7.18] and the fact that, as noted in Remark 3.12, the proof of Theorem 3.1 holds for $M_{\mathrm{Dol}}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)$. Furthermore, $\Gamma$ acts trivially on the cohomology of the fibers. This is because the action on the fibers can be extended to the multiplication action of the connected group $\mathrm{Hom}(\pi_{1}(S_{g,1}),\mathbb{C}^{*})\cong(\mathbb{C}^{*})^{2g}$ on itself. Then, since $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)\to\mathcal{M}_{g,1}$ is topologically locally trivial, we have $(R\pi_{\operatorname{GL}_{1}*}\mathbb{Q})^{\widetilde{\Gamma}}=R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}$. Thus, we conclude that